OpenStudy (anonymous):

y = x^2 - 3x - 8 What is the value of x if y = 0 ?

OpenStudy (yttrium):

subs y= 0 0 = x^2 - 3x - 8 by quadratic formula: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a } = \frac{ -(-3) \pm \sqrt{(-3)^1-4(1)(-8)} }{ (2)(1) }\] ^that would be tha values of x when y =0